OpenStudy (anonymous):

Given f(x) and g(x) below, find and simplify (f o g) (x). f(x)=(x)/(x+1) g(x)=6/x (f o g) (x)=?

OpenStudy (turingtest):

(f o g) (x)=(6/x)/[(6/x)+1]=6/(6+x)

OpenStudy (anonymous):

thanks ur awesome